Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
ruggi, casanueva, valentini - 12:27 Friday 09 April 2021 (51368) Print this report
PRCL loop instability at 11.6 Hz

This morning the instability at 11.6 Hz decided to come out again, so we had the opportunity to test the solution. A small portion of PR MIR correction has been sent to the marionette vertical actuation, passing through a resonant filter centerend at 11.6 Hz. Some coefficients, from -0.1 to -0.2 have been tested. A slow decreasing trend has been observed soon, but the higher value did not result to be better than the lower, because we started to see some 'beating', which was considered a bad thing. The final coefficient has been set at -0.17.

As we can see in the attached plot, the slow decreasing of the resonance went on constantly, up to the total disappearing. The engagement of the TX automatic alignment full bandwidth did not change the trend. Even if the tuning of the driving could be not perfect, the unstable parasitic loop should have been removed.

Images attached to this report
Search Help